# EUROPEAN ORGANIZATION FOR NUCLEAR RESEARCH (CERN)

### v1

@@
[Uncaptioned image] CERN-EP-2025-106 LHCb-PAPER-2024-042 3 June 2025
@@
The invariant-mass distribution for the coherent photoproduction of dipions in ultraperipheral PbPb collisions is measured using data, corresponding to an integrated luminosity of $224.6\pm 9.6\,\upmu\text{b}^{-1}$, collected by the LHCb experiment in 2018 at a nucleon-nucleon centre-of-mass energy $\sqrt{s_{\scriptscriptstyle\text{NN}}}=5.02\text{\,Te\kern-1.00006ptV}$. The dominant contribution is due to the $\rho^{0}$ meson but a consistent description across the full invariant-mass range requires accounting for the $\omega$ meson and introducing two resonances at masses of $1350\pm 20\text{\,Me\kern-1.00006ptV}$ and $1790\pm 20$ MeV with widths of about 300 MeV. The cross-section for each meson is measured differentially in twelve bins of rapidity from 2.05 to 4.90. Significant nuclear suppression is observed for the $\rho^{0}$ meson compared to expectations based on photoproduction on the proton.
Submitted to JHEP
@@ 1 Introduction
Photoproduction of vector mesons on nuclei probes nuclear structure and is sensitive to nuclear shadowing and saturation effects [1, 2, 3, 4]. The process can be studied in PbPb collisions at the LHC, when a photon is emitted from one lead ion, fluctuates into a virtual quark pair, and collides with the other target ion. The interaction with the target is mediated by colourless QCD propagators, Reggeons or Pomerons [5], and is shown diagrammatically in Fig. 1. Since the net colour flow in coherent photoproduction is zero, only a single vector meson is produced with a clear gap in rapidity relative to the outgoing ions. At small impact parameters, additional strong interactions between the ions are likely, thereby destroying the rapidity gap. However, by moving beyond the range of the strong force, only electromagnetic sources need to be considered. Experimentally, this is achieved in ultraperipheral collisions (UPC) [6], which produce the distinctive signature of a single reconstructed meson and no additional particle production, with the ions remaining intact.
@@
This paper presents a study of the dipion final state produced in PbPb UPC, by the LHCb collaboration. Although dominated by the ${{\rho}^{0}}$ meson, contributions are also present from the $\omega$ meson, nonresonant production, as well as possible excited vector mesons. Measurements of the cross-sections for coherent ${{\rho}^{0}}$ and $\omega$ meson production are performed. The measured ${{\rho}^{0}}$ cross-section, $\sigma_{{\rm PbPb}\rightarrow{\rm Pb}\,{{\rho}^{0}}\,{\rm Pb}}$, together with a calculation of the photon flux, then allows the photoproduction cross-section on the nucleus, $\sigma_{\gamma{\rm Pb}\rightarrow{{\rho}^{0}}{\rm Pb}}$, to be derived. A comparison of this result with photoproduction on the proton, $\sigma_{\gamma p\rightarrow{{\rho}^{0}}p}$, allows the nuclear suppression factor to be determined [3, 7]. Naïvely, the ion-to-proton photoproduction ratio should scale as $A^{4/3}$ where $A$ is the number of nucleons. However, since the ${{\rho}^{0}}$ meson is produced before it hits the target [8], it may undergo multiple interactions with the nucleons, which suppress the observable cross-section. This can be taken into account by the Gribov–Glauber mechanism [9, 10] giving predictions [11] where the suppression depends on the total $\rho$-nucleon cross-section. Suppression in excess of this would require additional mechanisms for nuclear shadowing [12] such as saturation, which is expected to be enhanced in nuclear collisions [4] and at low values of Bjorken-$x$, the fractional momentum of the partons in the nucleon. In this respect, the forward region covered by the LHCb experiment is of particular interest since in ${{\rho}^{0}}$-meson photoproduction, $x$ values down to $10^{-6}$ are probed.
@@
The $\rho$ meson is a broad resonance decaying almost exclusively to two pions. Since its discovery in the early 1960s, there have been difficulties in measuring its mass and width. In photoproduction, the resonance is also skewed, which is attributed to sizeable interference with the nonresonant production of pion pairs. A 1974 review by Spital and Yennie [13] summarises the situation, which is essentially unchanged today, in observing the resonance shape from reconstructing ${{\rho}^{0}}\rightarrow{{\pi}^{+}}{{\pi}^{-}}$ decays. The most precise measurements of the mass and width come from $e^{+}e^{-}\rightarrow{{\pi}^{+}}{{\pi}^{-}}$ data and $\tau$ decays, where discrepancies between these measurements seem to have been resolved [14]. The quoted values in the PDG [15] have very small uncertainties on the mass, $m_{{\rho}^{0}}=775.26\pm 0.23\text{\,Me\kern-1.00006ptV}$, [^1] Natural units are used throughout this paper. and width, $\Gamma_{{\rho}^{0}}=147.4\pm 0.8\text{\,Me\kern-1.00006ptV}$, but there are several assumptions and model dependencies involved in this determination. The situation is far from settled with a recent paper finding a value for the mass about 20 MeV lower [16]. Consistency in the $\rho$ parameters and the dipion spectrum measured in different production mechanisms is important, not just for an understanding of light-meson spectroscopy, but also because these feed into the determination of the hadronic corrections to the anomalous muon magnetic moment, $g-2$ [17].
@@
Although dominated by the ${{\rho}^{0}}$ meson and interference with nonresonantly-produced dipions, a precise measurement of the dipion spectrum allows other resonances to be identified. Interference between the ${{\rho}^{0}}$ and $\omega$ mesons, which was observed in AuAu collisions by the STAR collaboration [18], in $ep$ collisions by the H1 collaboration [19], and in $p$Pb collisions by the CMS collaboration [20], is seen allowing the differential cross-section for $\omega$ production in UPC to be measured. The dipion spectrum also indicates the presence of other particles, likely to be excited ${{\rho}^{0}}$ mesons.
Measurements of UPC ${{\rho}^{0}}$-meson production have been performed by the STAR collaboration [21, 22, 18] at nucleon-nucleon centre-of-mass energies, $\sqrt{s_{\rm NN}}$, of 62.4 and 200 GeV on gold nuclei, by the CMS collaboration in $p$Pb collisions at 5.02 TeV, while the ALICE collaboration performed measurements on xenon at 5.44 TeV [23], and lead at 2.76 and 5.02 TeV [24, 25, 26]. Measurements of UPC ${J\mskip-3.0mu/\mskip-2.0mu\psi}$-meson production have also been performed by the STAR collaboration [27] and at the LHC by the ALICE [28, 29, 30], CMS [31] and LHCb [32] collaborations, where the nuclear suppression factor is consistent with, or exceeds, what is expected in saturation models [7].
@@ 2 The LHCb detector, data collection and simulation
The LHCb detector [33, 34] is a single-arm forward spectrometer covering the pseudorapidity range $52<\eta<$ 5, originally designed for the precision study of particles containing $b$ or $c$ quarks. The detector includes a high-precision tracking system consisting of a silicon-strip vertex detector surrounding the collision region, a large-area silicon-strip detector located upstream of a dipole magnet with a bending power of about 4 T m, and three stations of silicon-strip detectors and straw drift tubes placed downstream of the magnet. The tracking system provides a measurement of the momentum, $p$, of charged particles with a relative uncertainty that varies from 0.5% at low momentum to 1.0% at 200 GeV. The minimum distance of a track to a primary PbPb collision vertex is measured with a resolution of (15+29/$p_{\mathrm{T}}$) $\mu$m, where $p_{\mathrm{T}}$ is in GeV. Different types of charged hadrons are distinguished using information from two ring-imaging Cherenkov (RICH) detectors. Photons, electrons and hadrons are identified by a calorimeter system consisting of scintillating-pad (SPD) and preshower detectors, an electromagnetic and a hadronic calorimeter. Muons are identified by a system composed of alternating layers of iron filters and multiwire proportional chambers. The pseudorapidity coverage is extended by forward and backward shower counters (HeRSCheL) consisting of five planes of scintillators with three planes at 114.0, 19.7 and 7.5 m upstream of the LHCb detector, and two planes downstream at 20.0 and 114.0 m. The HeRSCheL detector [35] significantly extends the acceptance for detecting particles from dissociated nuclei by covering the pseudorapidity range of $5\lesssim|\eta|\lesssim 10$, enhancing the classification of central exclusive production and UPC events. The online event selection is performed by a trigger, which consists of a hardware stage, based on information from the calorimeter and muon systems, followed by a software stage, which applies a full event reconstruction.
The data used in this analysis were collected in PbPb collisions in 2018 at a nucleon-nucleon centre-of-mass energy of 5.02 TeV, which corresponds to a total integrated luminosity of $224.6\pm 9.6\,\upmu\text{b}^{-1}$ [36]. The data were selected by a low-multiplicity trigger at the hardware stage that required between one and twenty SPD deposits, which has a signal efficiency above 98% over most of the dipion mass and rapidity range, since only two SPD deposits are expected for the signal. At the software trigger stage, one or more tracks were required to be reconstructed in detectors located both upstream and downstream of the magnet.
The use of simulation is limited to correcting for geometrical acceptance, measuring inefficiencies that amount to about $\sim 3\%$, and constraining the yields of small backgrounds whose shapes are determined in data. The production and decay of ${{\rho}^{0}}$ mesons is generated by the STARlight program [37], which is also used to investigate the two-photon production of muon and electron pairs. In addition, since the kinematics of the decay are fully determined once the mass, rapidity and transverse momentum of the parent meson have been specified along with the decay angle in the rest-frame of the meson, the ROOT [38] routine TGenPhaseSpace is used to study other particle decays. The interaction of the generated particles with the detector, and its response, are modelled using the Geant4 toolkit [39, 40] as described in Ref. [41].
@@ 3 Selection of coherent UPC-produced dipions
Simulation is used to determine a fiducial region in which the detector reconstruction is efficient. Both tracks must have $p_{\mathrm{T}}>100\text{\,Me\kern-1.00006ptV}$ and $52<\eta<$ 5, and the parent meson must have a rapidity, $y$, in the range $4.902.05<y<$ 4.90 and a mass above 400 MeV. The angle, $\theta^{*}$, between the outgoing positively charged pion and the outgoing lead ion, in the rest frame of the parent meson, must satisfy $|\cos\theta^{*}|$ below a value that varies from 0.7, between rapidities of 3 and 4, to 0.2 at the edges of the detector.
@@
The ${{K}^{0}_{\mathrm{S}}}\rightarrow{{\pi}^{+}}{{\pi}^{-}}$ sample in data is also used to determine the pion-identification efficiency, which is typically $(90\pm 3)\%$ where the uncertainty arises from kinematic differences between the $\rho$ and ${{K}^{0}_{\mathrm{S}}}$ mesons. The efficiency due to the requirement that there are no photons reconstructed in the event is calculated per rapidity bin, using events with identified photons but a well-balanced $p_{\mathrm{T}}$ between the two pions. This is typically 94% and a systematic uncertainty of 2% is assigned as the largest difference with simulation. The vetoes on identified electrons or muons have typical efficiencies of $(98\pm 1)\%$. The efficiency for the HeRSCheL veto on activity varies from 80% to 85% with rapidity. That veto is estimated to remove about half of the events having an additional electromagnetic interaction between the nuclei [42], which leads to excitation of one or both ions, a subsequent neutron emission (seen using zero-degree calorimeters by the ALICE and CMS [31, 26] collaborations), and partial or full breakup of the nucleus. Such events are identified by the presence of a diffractive peak in the $p_{\mathrm{T}}$ distribution of the vetoed events and a systematic uncertainty of 2% is assigned, reflecting the precision with which the shape of the incoherent component is determined. Efficiencies for the other selection requirements are above 94%. The total systematic uncertainty on the efficiency of the selection is about 7% of which 4% is uncorrelated between rapidity bins.
@@ 4 Fit to the dipion spectrum
where $\cal S$ is the signal function with free parameters $\alpha_{i}$, $\varepsilon$ is the efficiency and acceptance correction, and $R$ is a resolution function obtained from simulation and parametrised by the sum of two Gaussian functions. The narrower Gaussian is used to describe dipion candidates whose associated tracks are fully reconstructed by the tracking system and has a width that varies linearly from 3.5 to 8.0 MeV with rapidity and contributes a fraction that varies from 50% to 95% with rapidity. The wider Gaussian has a width of 50 MeV and accounts for candidates where one of the pions does not reach the tracking stations downstream of the magnet, usually due to low-momentum particles being swept out of the acceptance by the magnetic field. The background contributes less than 1% for $y>3.25$ and its shape, $\cal B$, is described using simulation and control samples in which the respective individual background source has been enhanced. For $y<3.25$ the background is only significant at the lowest dipion masses: above 600 MeV it is less than 3%.
For the signal function we start by considering the parametrisation of the dipion spectrum due to Söding [43], which models the $\rho$ meson using a Breit–Wigner function with mass, $m_{\rho}$, and an energy-dependent width, $\Gamma$, and includes interference with a nonresonant contribution described using a constant:
@@
Experiment LHCb ZEUS [44] H1 [45, 46] ALICE [24]
@@
The fit is performed for $m_{\pi\pi}$ between 400 MeV and 1200 MeV in twelve bins of rapidity. Compared to measurements at other experiments [44, 46, 45, 19, 21, 22, 18, 20, 24, 25], it is possible to extend the fit to lower mass values due to a reduced background contamination and the ability to reconstruct tracks down to a $p_{\mathrm{T}}$ of 100 MeV. The results of the fit for the rapidity range $4.502.75<y<$ 4.50 are given in Table 1 and are similar to those from other experiments for the mass and width, while the nonresonant contribution differs between experiments, most likely due to different amounts of residual background. Due to the large signal sample, the statistical uncertainties on the fit parameters are below 0.1%. A systematic uncertainty is estimated from the standard deviation of the results in the individual rapidity bins and by fixing the background contribution to zero.
@@
The fit is shown in Fig. 2. Although the overall shape is similar to the data, the fit quality is poor, particularly in the region of the $\omega$ resonance. Similar observations were made by the STAR [22], H1 [19], and CMS [20] collaborations and led to the introduction of a second Breit–Wigner function. Following the STAR parametrisation, the signal function is extended to read
@@
where $f_{\omega}$ is the contribution of the $\omega$ resonance and $\phi_{\omega}$ is its phase relative to the $\rho$ meson. The mass, $m_{\omega}$, and width, $\Gamma_{\omega}$, are fixed to the PDG values [15] and a constant width is assumed for $\Gamma_{\omega}$ as it is very narrow: implementing an energy-dependent width for the $\omega$ resonance makes minimal difference to the fit results. The left panel of Fig. 3 shows that the quality of the fit improves considerably in the peak region, and although the absolute size of the $\omega$ contribution is small, its presence and the resulting $\rho-\omega$ interference has a significant impact. The parameters of the fit are presented in Table 2 and are consistent with the STAR results.
@@
Although the fit is acceptable in the fitted region, when expanding the fit range up to dipion masses of 2.3 GeV, displayed in the left panel of Fig. 4, the model overshoots the data by an order of magnitude. Furthermore, the data show a clear structure at around 1.7 GeV, which is also discernible in STAR [47] and ALICE [25] data. To improve the quality of the fit to the extended mass range, two modifications are made to the signal model that now takes the form:
$\begin{split}{\cal S}(m_{\pi\pi})=&\frac{q^{3}(m_{\pi\pi})}{q^{3}(m_{\rho})}\Bigg{|}f_{\rho}\,{\cal{BW}}_{\rho}(m_{\pi\pi})\biggl{(}1+f_{\omega}\exp({i\phi_{\omega}})\frac{m_{\pi\pi}^{2}}{m_{\omega}^{2}}{\cal{BW}}_{\omega}(m_{\pi\pi})\biggr{)}\\ &+\frac{f_{\rm nr}\exp(i\phi_{\rm nr})}{(m_{\pi\pi}^{2}-4m_{\pi}^{2}+\Lambda^{2})^{\delta}}+f_{\rho\prime}\exp({i\phi_{\rho\prime}}){\cal{BW}}_{\rho\prime}(m_{\pi\pi})+f_{\rho\prime\prime}\exp({i\phi_{\rho\prime\prime}}){\cal{BW}}_{\rho\prime\prime}(m_{\pi\pi})\Bigg{|}^{2}\,.\\ \end{split}~{}$ (7)

### v2

@@
[Uncaptioned image] CERN-EP-2025-106 LHCb-PAPER-2024-042 18 November 2025
@@
The invariant-mass distribution for the coherent photoproduction of dipions in ultraperipheral PbPb collisions is measured using data, corresponding to an integrated luminosity of $224.6\pm 9.6\,\upmu\text{b}^{-1}$, collected by the LHCb experiment in 2018 at a nucleon-nucleon centre-of-mass energy $\sqrt{s_{\scriptscriptstyle\text{NN}}}=5.02\text{\,Te\kern-1.00006ptV}$. In the mass range from 400 to 1200 MeV, the results are consistent with previous experiments, with the spectrum dominated by the ${\rho}^{0}$} meson, which interferes with a nonresonant component, together with a smaller $\omega$ meson contribution. In an extended mass range up to 2300 MeV, models previously used do not fit the data and a consistent description requires the introduction of two resonances at masses of $1350\pm 20\text{\,Me\kern-1.00006ptV}$ and $1790\pm 20$ MeV with widths of about 300 MeV. The cross-section for each meson is measured differentially in twelve bins of rapidity from 2.05 to 4.90. The $\rho^{0}$ cross-section increases with rapidity from about 400 to 600 mb and is measured with a typical precision of 8%, while the cross-section times branching fraction for the $\omega,\rho^{\prime}$ and $\rho^{\prime\prime}$, with the statistical precision of the data, do not have a pronounced rapidity dependence and are between 0.5 and 1.5mb, with uncertainties up to 30%. A large nuclear suppression is observed for the $\rho^{0}$ meson compared to expectations based on photoproduction on the proton that use the impulse approximation. Significant suppression is also observed compared to that predicted by elastic scattering described in the Glauber approach, or with the addition of inelastic scattering in a Gribov–Glauber model.
Published in JHEP 11 (2025) 103.
@@ 1 Introduction
Photoproduction of vector mesons on nuclei probes nuclear structure and is sensitive to nuclear shadowing and saturation effects [Brodsky:1994kf, Klein:2019qfb, Guzey:2020ntc, Accardi:2012qut]. The process can be studied in PbPb collisions at the LHC, when a photon is emitted from one lead ion, fluctuates into a virtual quark pair, and collides with the other target ion. The interaction with the target is mediated by colourless QCD propagators, Reggeons or Pomerons [Donnachie:2002en], and is shown diagrammatically in Fig. 1. Since the net colour flow in coherent photoproduction is zero, only a single vector meson is produced with a clear gap in rapidity relative to the outgoing ions. At small impact parameters, additional strong interactions between the ions are likely, thereby destroying the rapidity gap. However, by moving beyond the range of the strong force, only electromagnetic sources need to be considered. Experimentally, this is achieved in ultraperipheral collisions (UPC) [Bertulani:2005ru], which produce the distinctive signature of a single reconstructed meson and no additional particle production, with the ions remaining intact.
@@
This paper presents a study of the dipion final state produced in PbPb UPC, by the LHCb collaboration. Although dominated by the ${\rho}^{0}$} meson, contributions are also present from the $\omega$ meson, nonresonant production, as well as possible excited vector mesons. Measurements of the cross-sections for coherent ${\rho}^{0}$} and $\omega$ meson production are performed. The measured ${\rho}^{0}$} cross-section, $\sigma_{{\rm PbPb}\rightarrow{\rm Pb}\,{{\rho}^{0}}\,{\rm Pb}}$, together with a calculation of the photon flux, then allows the photoproduction cross-section on the nucleus, $\sigma_{\gamma{\rm Pb}\rightarrow{{\rho}^{0}}{\rm Pb}}$, to be derived. A comparison of this result with photoproduction on the proton, $\sigma_{\gamma p\rightarrow{{\rho}^{0}}p}$, allows the nuclear suppression factor to be determined [Guzey:2020ntc, Mantysaari:2023xcu]. Naïvely, in the impulse approximation [Chew:1952fca], the ion-to-proton photoproduction ratio should scale as $A^{4/3}$ where $A$ is the number of nucleons. However, since the ${\rho}^{0}$} meson is produced before it hits the target [Ioffe:1969kf], it may undergo multiple interactions with the nucleons, which suppress the observable cross-section. This can be taken into account by the Gribov–Glauber mechanism [Gribov:1968jf, Glauber:1970jm] giving predictions [Guzey:2016piu] where the suppression depends on the total $\rho$-nucleon cross-section. The suppression for ${\rho}^{0}$} mesons is much larger than for $J\mskip-3.0mu/\mskip-2.0mu\psi$}} mesons, since the total ${\rho}^{0}$}–nucleon cross-section is about 20 mb, three order of magnitude greater than the $J\mskip-3.0mu/\mskip-2.0mu\psi$}-nucleon cross-section, which leads to a suppression of $\sigma_{\gamma{\rm Pb}\rightarrow{{\rho}^{0}}{\rm Pb}}$ compared to the impulse approximation by a factor five for the ${\rho}^{0}$, in contrast to about 10% for the $J\mskip-3.0mu/\mskip-2.0mu\psi$} meson. Suppression in excess of this would require additional mechanisms for nuclear shadowing [Guzey:2016qwo] such as saturation, which is expected to be enhanced in nuclear collisions [Accardi:2012qut] and at low values of Bjorken-$x$, the fractional momentum of the partons in the nucleon. In this respect, the forward region covered by the LHCb experiment is of particular interest since in ${\rho}^{0}$}-meson photoproduction, $x$ values down to $10^{-6}$ are probed.
@@
The $\rho$ meson is a broad resonance decaying almost exclusively to two pions. Since its discovery in the early 1960s, there have been difficulties in measuring its mass and width. In photoproduction, the resonance is also skewed, which is attributed to sizeable interference with the nonresonant production of pion pairs. A 1974 review by Spital and Yennie [Spital:1974cx] summarises the situation, which is essentially unchanged today, in observing the resonance shape from reconstructing ${{\rho}^{0}}\rightarrow{{\pi}^{+}}{{\pi}^{-}}$ decays. The most precise measurements of the mass and width come from $e^{+}e^{-}\rightarrow{{\pi}^{+}}{{\pi}^{-}}$ data and $\tau$ decays, where discrepancies between these measurements seem to have been resolved [Jegerlehner:2011ti]. The quoted values in the PDG [PDG2024] have very small uncertainties on the mass, $m_{{\rho}^{0}}=775.26\pm 0.23\text{\,Me\kern-1.00006ptV}$, [^1] Natural units are used throughout this paper. and width, $\Gamma_{{\rho}^{0}}=147.4\pm 0.8\text{\,Me\kern-1.00006ptV}$, but there are several assumptions and model dependencies involved in this determination. The situation is far from settled with a recent paper finding a value for the mass about 20 MeV lower [Bartos:2017ils]. Consistency in the $\rho$ parameters and the dipion spectrum measured in different production mechanisms is important, not just for an understanding of light-meson spectroscopy, but also because these feed into the determination of the hadronic corrections to the anomalous muon magnetic moment, g2g-2 [Davier:2017zfy].
@@
Although dominated by the ${\rho}^{0}$} meson and interference with nonresonantly-produced dipions, a precise measurement of the dipion spectrum allows other resonances to be identified. Interference between the ${\rho}^{0}$} and $\omega$ mesons, which was observed in AuAu collisions by the STAR collaboration [STAR:2017enh], in $ep$ collisions by the H1 collaboration [H1:2020lzc], and in ppPb collisions by the CMS collaboration [CMS:2019awk], is seen allowing the differential cross-section for $\omega$ production in UPC to be measured. The dipion spectrum also indicates the presence of other particles, likely to be excited ${\rho}^{0}$} mesons. In Ref. [ALICE:2020ugp], the ALICE collaboration reports an excess of events in the high-mass region, corresponding to a resonance mass of $1725\pm 17\text{\,Me\kern-1.00006ptV}$ and a width of $143\pm 21\text{\,Me\kern-1.00006ptV}$, as determined with a simple Gaussian model. With a similar model, the STAR collaboration [Klein:2016dtn] obtained a mass of $1653\pm 10\text{\,Me\kern-1.00006ptV}$ and a width of $164\pm 15\text{\,Me\kern-1.00006ptV}$, which they noted was compatible with the spin-3 $\rho_{3}(1690)$.
Measurements of UPC ${\rho}^{0}$}-meson production have been performed by the STAR collaboration [STAR:2007elq, STAR:2011wtm, STAR:2017enh] at nucleon-nucleon centre-of-mass energies, $\sqrt{s_{\rm NN}}$, of 62.4 and 200 GeV on gold nuclei, by the CMS collaboration in ppPb collisions at 5.02 TeV, while the ALICE collaboration performed measurements on xenon at 5.44 TeV [ALICE:2021jnv], and lead at 2.76 and 5.02 TeV [ALICE:2015nbw, ALICE:2020ugp, ALICE:2023jgu]. The STAR and ALICE data on coherent $\rho$ production in heavy-ion collisions were compared to the impulse approximation , Glauber, and Gribov–Glauber models in Ref.[Guzey:2016piu], where it was shown that the impulse approximation was about a factor four above the data, while the Glauber model was about a factor 1.5 too high. With the implementation of a Gribov–Glauber model, agreement with the data was achieved, although this used a parametrisation of H1 data that has since been improved [H1:2020lzc].
Recent discussion on nuclear suppression in coherent production has revolved around measurement of the $J\mskip-3.0mu/\mskip-2.0mu\psi$} meson, which have been performed by the STAR collaboration [STAR:2023vvb] and at the LHC by the ALICE [ALICE:2021tyx, ALICE:2021gpt, ALICE:2018oyo], CMS [CMS:2023snh] and LHCb [LHCb-PAPER-2022-012] collaborations, where the nuclear suppression factor is consistent with, or exceeds, what is expected in saturation models [Mantysaari:2023xcu]. Complementary measurements for the $\rho$ meson are welcome, since here the effect is much larger, although there are greater theoretical uncertainties when working in the nonperturbative regime.
@@ 2 The LHCb detector, data collection and simulation
[…]
The data used in this analysis were collected in PbPb collisions in 2018 at a nucleon-nucleon centre-of-mass energy of 5.02 TeV, which corresponds to a total integrated luminosity of $224.6\pm 9.6\,\upmu\text{b}^{-1}$ [LHCb:2014vhh]. The data were selected by a low-multiplicity trigger at the hardware stage that required between one and twenty SPD deposits, which has a signal efficiency above 98% over most of the dipion mass and rapidity range, since only two SPD deposits are expected for the signal. At the software trigger stage, one or more tracks were required to be reconstructed in detectors located both upstream and downstream of the magnet.
The use of simulation is limited to correcting for geometrical acceptance, measuring inefficiencies that amount to about $\sim 3\%$, and constraining the yields of small backgrounds whose shapes are determined in data. The production and decay of ${\rho}^{0}$} mesons is generated by the STARlight program [Klein:2016yzr], which is also used to investigate the two-photon production of muon and electron pairs. In addition, since the kinematics of the decay are fully determined once the mass, rapidity and transverse momentum of the parent meson have been specified along with the decay angle in the rest-frame of the meson, the ROOT [Antcheva:2011zz] routine TGenPhaseSpace is used to study other particle decays. The interaction of the generated particles with the detector, and its response, are modelled using the Geant4 toolkit [GEANT4:2002zbu, Allison:2006ve] as described in Ref. [Clemencic:2011zza].
@@ 3 Selection of coherent UPC-produced dipions
Simulation is used to determine a fiducial region in which the detector reconstruction is efficient. Both tracks must have $p_{\mathrm{T}}>100\text{\,Me\kern-1.00006ptV}$ and $2<\eta<5$, and the parent meson must have a rapidity, $y$, in the range $2.05<y<4.90$ and a mass above 400 MeV. The angle, $\theta^{*}$, between the outgoing positively charged pion and the outgoing lead ion, in the rest frame of the parent meson, must satisfy $|\cos\theta^{*}|$ below a value that varies from 0.7, between rapidities of 3 and 4, to 0.2 at the edges of the detector.
@@
The ${{K}^{0}_{\mathrm{S}}}\rightarrow{{\pi}^{+}}{{\pi}^{-}}$ sample in data is also used to determine the pion-identification efficiency, which is typically $(90\pm 3)\%$ where the uncertainty arises from kinematic differences between the $\rho$ and ${K}^{0}_{\mathrm{S}}$} mesons. The efficiency due to the requirement that there are no photons reconstructed in the event is calculated per rapidity bin, using events with identified photons but a well-balanced $p_{\mathrm{T}}$ between the two pions. This is typically 94% and a systematic uncertainty of 2% is assigned as the largest difference with simulation. The vetoes on identified electrons or muons have typical efficiencies of $(98\pm 1)\%$. The efficiency for the HeRSCheL veto on activity varies from 80% to 85% with rapidity. That veto is estimated to remove about half of the events having an additional electromagnetic interaction between the nuclei [Guzey:2013jaa], which leads to excitation of one or both ions, a subsequent neutron emission (seen using zero-degree calorimeters by the ALICE and CMS [CMS:2023snh, ALICE:2023jgu] collaborations), and partial or full breakup of the nucleus. Such events are identified by the presence of a diffractive peak in the $p_{\mathrm{T}}$ distribution of the vetoed events and a systematic uncertainty of 2% is assigned, reflecting the precision with which the shape of the incoherent component is determined. Efficiencies for the other selection requirements are above 94%. The total systematic uncertainty on the efficiency of the selection is about 7% of which 4% is uncorrelated between rapidity bins.
@@ 4 Fit to the dipion spectrum
[…]
For the signal function, several different parametrisations are used. First, a model, due to Söding [Soding:1965nh] that has contributions for a $\rho$ resonance and a nonresonant continuum that is constant with mass. Second, an extension to this that includes a contribution for the $\omega$ resonance. The third model uses a more flexible shape to describe the continuum. Finally, two additional resonances are introduced to describe the spectrum in an extended mass range.
The first parametrisation of the dipion spectrum is due to Söding [Soding:1965nh], which models the $\rho$ meson using a Breit–Wigner function with mass, $m_{\rho}$, and an energy-dependent width, $\Gamma$, and includes interference with a nonresonant contribution described using a constant:
@@
Experiment LHCb ZEUS [ZEUS:1997rof] H1 [H1:2009cml, H1:1996prv] ALICE [ALICE:2015nbw]
@@
The fit is performed for $m_{\pi\pi}$ between 400 MeV and 1200 MeV in twelve bins of rapidity. Compared to measurements at other experiments [ZEUS:1997rof, H1:1996prv, H1:2009cml, H1:2020lzc, STAR:2007elq, STAR:2011wtm, STAR:2017enh, CMS:2019awk, ALICE:2015nbw, ALICE:2020ugp], it is possible to extend the fit to lower mass values due to a reduced background contamination and the ability to reconstruct tracks down to a $p_{\mathrm{T}}$ of 100 MeV. The results of the fit for the rapidity range $2.75<y<4.50$ are given in Table 1 and are similar to those from other experiments for the mass and width, while the nonresonant contribution differs between experiments, most likely due to different amounts of residual background. Due to the large signal sample, the statistical uncertainties on the fit parameters are below 0.1%. A systematic uncertainty is estimated from the standard deviation of the results in the individual rapidity bins and by fixing the background contribution to zero.
@@
The fit is shown in Fig. 2. Although the overall shape is similar to the data, the fit quality is poor, particularly in the region of the $\omega$ resonance. Similar observations were made by the STAR [STAR:2011wtm], H1 [H1:2020lzc], and CMS [CMS:2019awk] collaborations and led to the introduction of a second Breit–Wigner function. Following the STAR parametrisation, the signal function is extended to read
@@
where $f_{\omega}$ is the contribution of the $\omega$ resonance and $\phi_{\omega}$ is its phase relative to the $\rho$ meson. The mass, $m_{\omega}$, and width, $\Gamma_{\omega}$, are fixed to the PDG values [PDG2024] and a constant width is assumed for $\Gamma_{\omega}$ as it is very narrow: implementing an energy-dependent width for the $\omega$ resonance makes minimal difference to the fit results. The left panel of Fig. 3 shows that the quality of the fit improves considerably in the peak region, and although the absolute size of the $\omega$ contribution is small, its presence and the resulting $\rho-\omega$ interference has a significant impact. The parameters of the fit are presented in Table 2 and are consistent with the STAR results.
@@
Although the fit is acceptable in the fitted region, when expanding the fit range up to dipion masses of 2.3 GeV, displayed in the left panel of Fig. 4, the model overshoots the data by an order of magnitude. Furthermore, the data show a clear structure at around 1.7 GeV, which is also discernible in STAR [Klein:2016dtn] and ALICE [ALICE:2020ugp] data. To improve the quality of the fit to the extended mass range, two modifications are made to the signal model that now takes the form:
[…]